OpenStudy (anonymous):

(sqrt of 2)^(x+4)=4^x How do I get the bases to be the same?

myininaya (myininaya):

4=(sqrt(2))^4

myininaya (myininaya):

(sqrt(2))^(x+4)=(sqrt(2))^(4x)

myininaya (myininaya):

so x+4=4x

OpenStudy (anonymous):

I don't understand what you did there.

myininaya (myininaya):

since the bases are the same, the exponents have to be the same

OpenStudy (anonymous):

Oh, you just simplified the 4 into sqrt of 2^4 before including the exponents?

myininaya (myininaya):

yes becauce sqrt(2)*sqrt(2)*sqrt(2)*sqrt(2)=4

myininaya (myininaya):

so both sides would have the same base

OpenStudy (anonymous):

Okay great, that I understand :D I can do the rest, thank you very much :D

myininaya (myininaya):

you're welcome

OpenStudy (anonymous):

so x=4/3 then?

OpenStudy (anonymous):

Yes, that is right :D

myininaya (myininaya):

correct!